# Assessing Large Language Models in Updating Their Forecasts with New Information

### v1

@@ 5.1 Models
We evaluate three openly available reasoning models from the DeepSeek-R1 series (Guo et al., 2025): DeepSeek-R1-Distill-Qwen-1.5B, DeepSeek-R1-Distill-Qwen-7B, and DeepSeek-R1-Distill-LLaMA-8B, together with their published base counterparts from the same backbones (Qwen 2.5 (Bai et al., 2025) and Llama 3.1 (Dubey et al., 2024) families, as indicated in the official model cards). [^2] We pair each R1-distilled model with the corresponding base instruction model from its backbone family following the mapping described in the providers’ model cards. This choice serves two purposes. First, it provides coverage across both Qwen and Llama backbones and multiple parameter scales (1.5B/7B/8B), enabling us to probe how scale and architecture relate to belief-updating behavior. Second and critically, it allows us to isolate the effect of reasoning-style post-training (e.g., distillation/RL procedures specific to R1) by comparing each R1 model against a matched base model. Including unrelated weaker models would introduce confounds (different tokenizers, pretraining corpora, and objectives) that obscure the contribution of reasoning post-training and inflate variance without strengthening causal claims about update dynamics. All models are evaluated with identical prompts and decoding settings.
@@ 5.3 Confidence Extraction
The model is instructed to provide a binary prediction (“Yes” or “No”) and to assign a probability on a 1–10 scale with descriptive anchors (“1: extremely unlikely” $\ldots$ “10: extremely likely”). We normalize this verbalized score to $[0,1]$ for evaluation. This setting aligns with how human forecasters typically express uncertainty through explicit probability estimates, facilitating direct comparison with the aggregated human reference.
@@ 6.2 Ablation: Accumulated News Context
Table 2: Ablation on news context. Models are evaluated with either only the most recent update (S = Single) or the full sequence of updates from $T_{0}$ to $T_{1}$ (A = Accumulated). Metrics are grouped as in Table 1(b). Results for magnitude error and calibration change will be added once available.
@@
(a) DS R1 Qwen-1.5B (S)
(b) DS R1 Qwen-7B (S)
(c) DS R1 LLaMA-8B (S)
(d) DS R1 Qwen-1.5B (A)
(e) DS R1 Qwen-7B (A)
(f) DS R1 LLaMA-8B (A)
@@ 6.3 Ablation: Directional QA Prompting
Table 3 reports directional agreement metrics under this setting. Compared to the probability-based approach, the direct method yields markedly higher MDA and F1, especially for larger models. For example, DS R1 LLaMA-8B achieves nearly 0.49 MDA with single updates, compared to only 0.34 under verbalized confidence (cf. Table 1(b)). This indicates that models are substantially better at reasoning about the qualitative direction of change than at producing well-calibrated probabilities. Similar as before, accumulated news again does not improve performance: both RS R1 Qwen-7B and RS R1 LLaMA-8B see declines in MDA and F1 when given the full sequence of updates. Together with Sec. 6.2, this highlights that concatenating news does not help models integrate evidence and may in fact dilute the signal. Overall, this ablation suggests that while LLMs struggle with precise probability calibration, they are relatively more reliable at qualitative directional reasoning.
@@ 7 Conclusion
We introduced EvolveCast, the first framework to assess how language models revise forecasts when new evidence emerges. Across multiple models and extensive ablation studies, we find that updates are often conservative or inconsistent, with neither verbalized nor logit-based confidence clearly superior and both far from human references. These results underscore the challenge of belief updating in current LLMs and the need for more robust approaches to handling evolving evidence.
Ethics Statement
@@
Reproducibility Statement
@@
Acknowledgements
All authors of this paper are supported by the ERC grant AVeriTeC (GA 865958). Andreas Vlachos receives further support from the DARPA program SciFy.
Appendix A Limitations
Our study is limited in scope in several ways. We focus on binary forecasting questions, which provides clarity for evaluation but does not cover all types of forecasting tasks. We also evaluate a small set of openly available reasoning models, so results should not be overgeneralized. Finally, because the data construction relies on public APIs and human discussion timestamps, exact replication of forecaster information exposure is not possible. These constraints are inherent to working with real-world forecasting data, and we have aimed to minimize their impact in this paper.
Appendix B The Use of Large Language Models
@@ G.1 Similarity-Sensitive Confidence
In this ablation, we implement a semantics-aware confidence estimation method inspired by clustering-based approaches (Kuhn et al., 2023). The intuition is that if a model generates multiple plausible answers, the distribution of these generations can provide a more robust uncertainty estimate than any single output. Concretely, we sample $N$ candidate outputs $\{y^{(1)},\dots,y^{(N)}\}$ and group them into $K$ clusters based on cosine similarity of sentence embeddings (Reimers & Gurevych, 2019). Confidence for cluster $C_{k}$ is then defined as

### v2

@@ 5.1 Models
We evaluate three openly available reasoning models from the DeepSeek-R1 series (Guo et al., 2025): DeepSeek-R1-Distill-Qwen-1.5B and 7B, and DeepSeek-R1-Distill-LLaMA-8B, together with their published base counterparts from the same backbones (Qwen 2.5 (Bai et al., 2025) and Llama 3.1 (Dubey et al., 2024) families, as indicated in the official model cards). [^2] We pair each R1-distilled model with the corresponding base instruction model from its backbone family following the mapping described in the providers’ model cards. This choice serves two purposes. First, it provides coverage across both Qwen and Llama backbones and multiple parameter scales (1.5B/7B/8B), enabling us to probe how scale and architecture relate to belief-updating behavior. Second and critically, it allows us to isolate the effect of reasoning-style post-training (e.g., distillation/RL procedures specific to R1) by comparing each R1 model against a matched base model. Including unrelated weaker models would introduce confounds (different tokenizers, pretraining corpora, and objectives) that obscure the contribution of reasoning post-training and inflate variance without strengthening causal claims about update dynamics. All models are evaluated with identical prompts and decoding settings.
@@ 5.3 Confidence Extraction
The model is instructed to provide a binary prediction (“Yes” or “No”) and to assign a probability on a 1–10 scale with descriptive anchors (“1: extremely unlikely” $\ldots$ “10: extremely likely”). We normalize this verbalized score to $[0,1]$. This setting aligns with how human forecasters typically express uncertainty through explicit probability estimates, facilitating direct comparison with the aggregated human reference.
@@
6 Results
This consistent conservatism likely stems from the tension between the model’s pre-trained world model (parametric memory) and the external context. Unlike a Bayesian agent that updates its prior based on the likelihood of new evidence, the LLM processes the retrieved news merely as additional tokens within the attention mechanism. Consequently, the massive weight of pre-training data acts as an overwhelming anchor, preventing the short context window from shifting the probability distribution as significantly as a rational update would require.
@@ 6.2 Ablation: Accumulated News Context
Table 2: Ablation on news context. Models are evaluated with either only the most recent update (S = Single) or the full sequence of updates from $T_{0}$ to $T_{1}$ (A = Accumulated). Metrics are grouped as in Table 1(b).
@@
(a) Qwen-1.5B (S)
(b) Qwen-7B (S)
(c) LLaMA-8B (S)
(d) Qwen-1.5B (A)
(e) Qwen-7B (A)
(f) LLaMA-8B (A)
@@
This pattern suggests that models struggle to hierarchically weight evidence, often failing to prioritize the most recent signals over older context. Instead of converging on a more precise estimate, the added history appears to introduce volatility, resulting in spurious directional shifts. This highlights a distinctive challenge for LLMs: unlike human forecasters who discount outdated news, models may treat accumulated evidence as competing noise.
@@ 6.3 Ablation: Directional QA Prompting
Table 3 reports directional agreement metrics under this setting. Compared to the probability-based approach, the direct method yields markedly higher MDA and F1, especially for larger models. For example, DS R1 LLaMA-8B achieves nearly 0.49 MDA with single updates, compared to only 0.34 under verbalized confidence (cf. Table 1(b)). This indicates that models are substantially better at reasoning about the qualitative direction of change than at producing well-calibrated probabilities. Similar as before, accumulated news again does not improve performance: both DS R1 Qwen-7B and DS R1 LLaMA-8B see declines in MDA and F1 when given the full sequence of updates. Together with Sec. 6.2, this highlights that concatenating news does not help models integrate evidence and may in fact dilute the signal. Overall, this ablation suggests that while LLMs struggle with precise probability calibration, they are relatively more reliable at qualitative directional reasoning.
Crucially, the higher performance in this directional setting serves as a validation of the benchmark itself: it confirms that the retrieved news snippets do contain accessible, relevant signals for updating beliefs. The underperformance in the main probabilistic setting is therefore not due to a lack of evidence in the text, but rather stems from the models’ inability to translate that evidence into calibrated shifts.
@@ 7 Conclusion
We introduced EvolveCast, the first framework to assess how language models revise forecasts when new evidence emerges. Across multiple models and ablation studies, we find that updates are often conservative or inconsistent, with neither verbalized nor logit-based confidence clearly superior and both far from human references. These results underscore the challenge of belief updating in current LLMs: simply retrieving new information is insufficient for rational forecasting and future work must go beyond static context.
Impact Statement
@@
Appendix A The Use of Large Language Models
Appendix B Reproducibility Statement
@@ G.1 Similarity-Sensitive Confidence
In this ablation, we implement a semantics-aware confidence estimation method inspired by clustering-based approaches (Kuhn et al., 2023). The intuition is that if a model generates multiple plausible answers, the distribution of these generations can provide a more robust uncertainty estimate than any single output. Concretely, we sample $N$ candidate outputs $\{y^{(1)},\dots,y^{(N)}\}$ and group them into $K$ clusters based on cosine similarity of sentence embeddings (Reimers and Gurevych, 2019). Confidence for cluster $C_{k}$ is then defined as